# Spectral Analysis and Decay Mechanisms of $1^{-+}$ Hybrid States in Light Meson Sector

### v2

@@ IV.2 The ground states of $I^{G}J^{PC}=0^{+}1^{-+}$
As mentioned in Section II, the mass of the pure $s\bar{s}g$ state is 2023 MeV, which is higher than 1855 MeV, whereas the pure $(u\bar{u}+d\bar{d})g/\sqrt{2}$ state has a mass lower than 1855 MeV. This suggests that $\eta_{1}(1855)$ is likely a mixed state of $s\bar{s}g$ and $(u\bar{u}+d\bar{d})g/\sqrt{2}$, as proposed by Swanson:2023zlm . This hypothesis also lends support to efforts to construct the $1^{-+}$ nonet Chen:2023ukh ; Qiu:2022ktc ; Eshraim:2020ucw .
@@
Similar to the flavor mixing treatment of meson decay width in Ref. Li:2008mza , the partial widths of $\eta_{1}^{H}$ and $\eta_{1}^{L}$ become with mixing:
$\Gamma(\eta_{1}^{H}\rightarrow BC)=G\sum_{LS}|\sin\phi{\cal{M}}^{LS}_{n\bar{n}\rightarrow BC}+\cos\phi{\cal{M}}^{LS}_{s\bar{s}\rightarrow BC}|^{2},$
$\Gamma(\eta_{1}^{L}\rightarrow BC)=G^{{}^{\prime}}\sum_{LS}|\cos\phi{\cal{M}}^{LS}_{n\bar{n}\rightarrow BC}-\sin\phi{\cal{M}}^{LS}_{s\bar{s}\rightarrow BC}|^{2}.$
Where $G=\frac{\alpha_{s}PE_{B}E_{C}}{\pi M_{\eta_{1}^{H}}}$, $G^{{}^{\prime}}=\frac{\alpha_{s}PE_{B}E_{C}}{\pi M_{\eta_{1}^{L}}}$, for the decay channels associated with $\eta$ and $\eta^{\prime}$, we regard them as
$\begin{pmatrix}\eta\\ \eta^{\prime}\end{pmatrix}$ $=\begin{pmatrix}\cos\theta_{i}&-\sin\theta_{i}\\ \sin\theta_{i}&\cos\theta_{i}\end{pmatrix}\begin{pmatrix}|n\bar{n}\rangle\\ |s\bar{s}\rangle\end{pmatrix},$ (34)
$\theta_{i}\approx 43.3^{\circ}$, see Ref. Wang:2020due .
In contrast, assuming that $\eta_{1}(1855)$ is a higher isospin partner, it exhibits properties that strongly support its interpretation as a mixed hybrid state. Table 6 presents the decay results. Its dominant decay channels, $a_{1}(1260)\pi$ and $K_{1}(1270)K$ align well with theoretical predictions (see Fig. 3(a)), and the mixing angle between $(u\bar{u}+d\bar{d})g/\sqrt{2}$ and $s\bar{s}g$ is constrained to the range of $21^{\circ}$ to $74^{\circ}$. This result presents a compelling argument for classifying $\eta_{1}(1855)$ as a hybrid meson and highlights the importance of mixing effects in understanding its decay dynamics.
For quantum numbers $I^{G}J^{PC}=0^{+}1^{-+}$, the state $(u\bar{u}+d\bar{d})g/\sqrt{2}$ is assigned as the lower isoscalar partner to the $\eta_{1}(1855)$ state, with the assumption that $\eta_{1}(1855)$ is the higher isoscalar partner. The lower state is labeled as $\eta_{1}(1600)$, with the mixing angle ranging from $21^{\circ}$ to $74^{\circ}$, as previously discussed. The results presented in Table 7 show that $a_{1}(1260)\pi$ is the characteristic decay channel. The total decay width of $\eta_{1}(1600)$, as a function of the mixing angle, is shown in Figure 3 (b). The decay width of the $\eta_{1}(1600)$ state is predicted to be narrow, consistent with results reported in Chen:2023ukh ; Shastry:2023ths .
The results highlight the need for precise experimental measurements of decay branching ratios, particularly for $\alpha_{s}$ of its decay channels will further constrain theoretical models and provide clarity on its hybrid nature. Similarly, targeted searches for the predicted excited states will be crucial in expanding our understanding of the $1^{-+}$ meson spectrum.
@@ IV.3 The first excited states of $I^{G}J^{PC}=0^{+}1^{-+}$
In accordance with the mass spectrum calculated in Section II, the first excitation state of $\eta_{1}(1855)$ is labeled as $\eta_{1}(2355)$, maintaining the same mixing angle as that of $\eta_{1}(1855)$. The decay width results are presented in Table 8. For a clearer visualization of the decay widths, refer to Fig. 4(a). The total decay width of $\eta_{1}(2355)$ is approximately 100 MeV, with dominant decay channels including $K_{1}(1270)K$, $K_{1}(1650)K$, and $a_{1}(1640)\pi$.
Additionally, the first excited state of $\eta_{1}(1600)$ is named $\eta_{1}(2193)$, serving as the partner of $\eta_{1}(2355)$. The decay widths are presented in Table 9. As shown in Fig. 4 (b), the main decay modes of $\eta_{1}(2193)$ include $K_{1}(1270)K$, $K_{1}(1650)K$, and $a_{1}(1640)\pi$. We anticipate that future experiments will investigate these channels to detect $\eta_{1}(2355)$ and $\eta_{1}(2193)$.
@@
$b_{1}(1235)\rho$ $a_{1}(1260)\pi$ $f_{1}(1420)\eta$
$1.08s^{2}$ $5.64s^{2}$ $\approx 0$
$K_{1}(1270)K$ $K_{1}(1400)K$ $K_{1}(1270)K^{*}$
$1s^{2}+17.6c^{2}+8.4cs$ $\approx 0$ $\approx 0$
$K_{1}(1650)K$ $\pi(1300)\pi$ $\pi_{2}(1670)\pi$
$22.84s^{2}+36.2c^{2}+57.5cs$ $6.3s^{2}$ $7.8s^{2}$
$K(1460)K$ $K^{*}(1410)K$ $a_{1}(1640)\pi$
$0.5s^{2}+3.36c^{2}+2.53cs$ $0.31s^{2}+2.08c^{2}+1.6cs$ $35.28s^{2}$
@@ V Discussion and conclusion
This study investigates the properties of hybrid $J^{PC}=1^{-+}$ mesons and their isoscalar counterparts through comprehensive mass and decay analyzes. Using a potential model inspired by the lattice QCD and the constituent-gluon framework, we calculate the masses and decay widths of key exotic mesons, shedding light on their structure and classification.
Our findings suggest that $\eta_{1}(1855)$ is a strong candidate for a mixed hybrid meson, composed of $(u\bar{u}+d\bar{d})g/\sqrt{2}$ and $s\bar{s}g$ components, with mixing angles constrained between $21^{\circ}$ to $74^{\circ}$. Its mass and decay properties, including dominant channels such as $a_{1}(1260)\pi$ and $K_{1}(1270)K$ , align well with theoretical predictions. These results support its classification as a hybrid meson within the meson nonet.
In contrast, the narrow decay widths predicted for $\pi_{1}(1600)$ and $\pi_{1}(2015)$ deviate significantly from experimental observations, challenging their classification as hybrid mesons. Alternative explanations, such as tetraquark or molecular configurations, merit further investigation. Our results also predict narrow width excited states, $\eta_{1}(2355)$ and $\eta_{1}(2193)$, which primarily decay into channels like $K_{1}(1270)K$ and $a_{1}(1640)\pi$. While challenging to detect experimentally, these states offer promising opportunities to test and validate hybrid meson models.
This work underscores the importance of precise experimental measurements to resolve ambiguities in hybrid meson classification. Future experiments should focus on measuring branching ratios for key decay channels, such as those of $\eta_{1}(1855)$, to refine mixing angle estimates. Furthermore, searches for the predicted excited states will be instrumental in advancing our understanding of hybrid meson dynamics and their role in exotic hadronic states.
The results presented here highlight the intricate interplay between theory and experiment in the study of exotic mesons. Ongoing collaboration between experimental efforts and theoretical advancements, including refined lattice QCD calculations and improved decay models, will be critical in uncovering the true nature of hybrid mesons and enhancing our understanding of the strong interaction.
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China under Grant No. 12405104, the Natural Science Foundation of Hebei Province under Grant No. A2022203026, the Higher Education Science and Technology Program of Hebei Province under Contract No. BJK2024176, and the Research and Cultivation Project of Yanshan University under Contract No. 2023LGQN010.

### v3

@@ IV.2 The ground states of $I^{G}J^{PC}=0^{+}1^{-+}$
As discussed in Section II, the mass of the pure $s\bar{s}g$ state is 2023 MeV, which exceeds 1855 MeV, while the pure $(u\bar{u}+d\bar{d})g/\sqrt{2}$ state has a mass below 1855 MeV. This indicates that $\eta_{1}(1855)$ is likely a mixture of $s\bar{s}g$ and $(u\bar{u}+d\bar{d})g/\sqrt{2}$ states, as proposed by Swanson:2023zlm . This hypothesis also supports efforts to construct the $1^{-+}$ nonet, as explored in Chen:2023ukh ; Qiu:2022ktc ; Eshraim:2020ucw .
@@
In contrast, if $\eta_{1}(1855)$ is interpreted as a higher isospin partner, its properties strongly suggest a mixed hybrid state. Table 6 summarizes its decay results. The dominant decay channels, $a_{1}(1260)\pi$ and $K_{1}(1270)\bar{K}$, align closely with theoretical predictions (see Fig. 3(a)), and the mixing angle between $(u\bar{u}+d\bar{d})g/\sqrt{2}$ and $s\bar{s}g$ is constrained to the range of $17.7^{\circ}$ to $84.2^{\circ}$. These findings provide compelling evidence for classifying $\eta_{1}(1855)$ as a hybrid meson and underscore the significance of mixing effects in its decay dynamics.
For quantum numbers $I^{G}J^{PC}=0^{+}1^{-+}$, the state $(u\bar{u}+d\bar{d})g/\sqrt{2}$ is assigned as the lower isoscalar partner to the $\eta_{1}(1855)$ state, with the assumption that $\eta_{1}(1855)$ is the higher isoscalar partner. The lower state is labeled as $\eta_{1}(1600)$, with the mixing angle ranging from $17.7^{\circ}$ to $84.2^{\circ}$, as previously discussed. The results presented in Table 7 show that $a_{1}(1260)\pi$ is the characteristic decay channel. The total decay width of $\eta_{1}(1600)$, as a function of the mixing angle, is shown in Figure 3 (b).
The results emphasize the necessity for precise experimental measurements of decay branching ratios, especially for $\alpha_{s}$. Such measurements will further constrain theoretical models and elucidate the hybrid nature of the particle. Additionally, targeted searches for the predicted excited states are essential for advancing our understanding of the $1^{-+}$ hybrid meson spectrum.
@@ IV.3 The first excited states of $I^{G}J^{PC}=0^{+}1^{-+}$
In accordance with the mass spectrum calculated in Section II, the first excitation state of $\eta_{1}(1855)$ is labeled as $\eta_{1}(2355)$, maintaining the same mixing angle as that of $\eta_{1}(1855)$. The decay width results are presented in Table 8. For a clearer visualization of the decay widths, refer to Fig. 4(a). The total decay width of $\eta_{1}(2355)$ is approximately 150 MeV, with dominant decay channels including $K_{1}(1270)\bar{K}$, $K_{1}(1650)\bar{K}$, and $a_{1}(1640)\pi$.
Additionally, the first excited state of $\eta_{1}(1600)$ is named $\eta_{1}(2193)$, serving as the partner of $\eta_{1}(2355)$. The decay widths are presented in Table 9. As shown in Fig. 4 (b), the main decay modes of $\eta_{1}(2193)$ include $K_{1}(1270)\bar{K}$, $K_{1}(1650)\bar{K}$, and $a_{1}(1640)\pi$. We anticipate that future experiments will investigate these channels to detect $\eta_{1}(2355)$ and $\eta_{1}(2193)$.
@@
$\eta(1295)\eta$ $\eta(1475)\eta$ $\pi(1300)\pi$
$0.49s^{2}+0.77c^{2}+1.23cs$ $\approx 0$ $8.4s^{2}$
$K(1460)\bar{K}$ $K^{*}(1410)\bar{K}$ $a_{1}(1260)\pi$
$0.49s^{2}+3.4c^{2}+2.6cs$ $0.31s^{2}+2.1c^{2}+1.6cs$ $7.5s^{2}$
$f_{1}(1285)\eta$ $f_{1}(1420)\eta$ $a_{1}(1640)\pi$
$\approx 0$ $\approx 0$ $44s^{2}$
$K_{1}(1270)\bar{K}$ $K_{1}(1400)\bar{K}$ $h_{1}(1170)\omega$
$1s^{2}+17.7c^{2}+8.4cs$ $\approx 0$ $0.75s^{2}$
$K_{1}(1650)\bar{K}$ $K_{2}(1770)\bar{K}$ $b_{1}(1235)\rho$
$46s^{2}+75c^{2}+117.4cs$ $\approx 0$ $1.5s^{2}$
$\eta_{2}(1645)\eta$ $\pi_{2}(1670)\pi$
$0.53s^{2}$ $11s^{2}$
@@ V Discussion and conclusion
This study investigates the properties of $J^{PC}=1^{-+}$ hybrid mesons and their isoscalar counterparts through comprehensive mass and decay analyzes. Using a potential model inspired by the lattice QCD and the constituent-gluon framework, we calculate the masses and decay widths of key exotic mesons, shedding light on their structure and classification.
Our findings indicate that $\eta_{1}(1855)$ is a strong candidate for a mixed hybrid meson, characterized by $(u\bar{u}+d\bar{d})g/\sqrt{2}$ and $s\bar{s}g$ components, with mixing angles constrained to the range of $17.7^{\circ}$ to $84.2^{\circ}$. Its mass and decay properties, particularly the dominant channels $a_{1}(1260)\pi$ and $K_{1}(1270)\bar{K}$, closely align with theoretical predictions. These results strongly support the classification of $\eta_{1}(1855)$ as a hybrid meson.
In contrast, the narrow decay widths predicted for $\pi_{1}(1600)$ and $\pi_{1}(2015)$ deviate significantly from experimental observations, challenging their classification as hybrid mesons. Alternative explanations, such as tetraquark or molecular configurations, merit further investigation. This study also predicts the excited states of $\eta_{1}(1855)$ and $\eta_{1}(1600)$, specifically $\eta_{1}(12355)$ and $\eta_{1}(2193)$, respectively. Their primary decay channels are identified as $K_{1}(1650)\bar{K}$, $K_{1}(1270)\bar{K}$, and $a_{1}(1640)\pi$. However, the relatively narrow width of $\eta_{1}(2193)$ presents a challenge for experimental detection.
This work emphasize the importance of precise experimental measurements to resolve ambiguities in hybrid meson classification. Future experiments should focus on measuring branching ratios for key decay channels, such as those of $\eta_{1}(1855)$, to refine mixing angle estimates. Furthermore, searches for the predicted excited states will be instrumental in advancing our understanding of hybrid meson dynamics and their role in exotic hadronic states.
The results presented in this study demonstrate the complex relationship between theoretical predictions and experimental investigations in the exploration of exotic mesons. Continued collaboration between experimental efforts and theoretical advancements, including refined lattice QCD calculations and improved decay models, will be critical in uncovering the true nature of hybrid mesons and enhancing our understanding of the strong interaction.
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China under Grant No. 12405104 and No. 12305087, the Start-up Funds of Nanjing Normal University under Grant No. 184080H201B20, the Natural Science Foundation of Hebei Province under Grant No. A2022203026, the Higher Education Science and Technology Program of Hebei Province under Contract No. BJK2024176, and the Research and Cultivation Project of Yanshan University under Contract No. 2023LGQN010.